# Comparing next-generation detector configurations for high-redshift gravitational wave sources with neural posterior estimation

### v1

@@ II.3 Metrics to assess detector performance
which quantifies the extent to which the prior volume has been constrained by the data–by a factor $2^{I}$ (Skilling:2004pqw). As opposed to single- or few-parameter metrics, the information gain fully characterizes the size of the 11-dimensional posterior distribution, accounting for correlations across any parameters. Equation 3 is evaluated using a Monte Carlo integral. Additional details on this are provided in Appendix E.
@@
Because Dingo-IS provides full parameter estimates for each source, we can define precise metrics to assess how different detector networks perform in source localization. For each event, we compute the area (in deg2) enclosed by the smallest 90% confidence region. The sky maps are sampled using equal-area HEALPix (Hierarchical Equal Area isoLatitude Pixelization; Gorski:1999rt; Gorski:2004by) pixels, where each of the $N$ pixels carries a posterior probability $p_{i}$ that the source lies within that pixel.
@@
We employ the ligo-skymap library (Singer:2015ema; Singer:2016eax; Singer:2016erz), which also identifies the number of disconnected sky modes (or probability islands) within a given area using a flood-fill algorithm (foley1996computer), which finds all pixels connected to a starting one. Additionally, by including the posterior on luminosity distance, we compute the 90% highest probability density comoving volume ($\Delta V^{c}_{90\%}$, see Section 3 of Singer:2016eax).
@@
The other reported interval widths, for example $\Delta x/x^{\mathrm{inj}}$, indicate the relative deviation of a parameter from its injected value. Here, $\Delta x$ corresponds to the square root of the associated covariance matrix element, representing the $1\sigma$ uncertainty in the estimate.
III Results
III.1 High signal-to-noise ratio vs. high sample efficiency
@@ III Results
Figure 1: Top panel: Injections as function of sample efficiency ($\epsilon$, $x$-axis) and optimal SNR ($\rho$, $y$-axis) for each detector configuration (color-coded). Filled markers indicate injections with sample efficiencies $1\%$. Percentages in parentheses denote the fraction of sources with sample efficiency above $1\%$, while the horizontal colored lines indicate the median $\rho$ for sources exceeding this threshold. Bottom panel: Injections as function of detector-frame chirp mass ($x$-axis) and luminosity distance (left $y$-axis), with corresponding redshifts (right $y$-axis). Colored markers indicate events with sample efficiency greater than $1\%$ in all considered detector configurations. See Section III.1 for details.
We generate $1000$ sets of BBH merger parameters sampled from the prior and use the same samples for all configurations.
@@
From the Dingo proposal distribution $q_{\phi}(\theta|d)$, we draw $10^{5}$ posterior samples and compute the corresponding importance weights (see Section II.2). For each injected signal, we then evaluate the sample efficiency. To ensure unbiased estimates of the inferred parameters, we require at least $10^{3}$ effective samples per source, corresponding to a sample efficiency >1>1 %. The top panel of Figure 1 shows that a large fraction of injected signals achieves a sample efficiency greater than $1\%$, ranging from $96\%$ for $\Delta$ configuration to a minimum of $85\%$ for 2L MisA + CE. These values remain consistently high, with even better performance for the $\Delta$ configuration than previously reported in Santoliquido:2025lot. This improvement arises from the increased dimensionality of the embedding network (see Section II.2), emphasizing the crucial role of this component within the NPE framework.
@@
The top panel of Figure 1 also shows that $50\%$ of the events have an optimal signal-to-noise ratio (SNR, see Appendix D) exceeding $40$ for the $\Delta$ configuration and up to $70$ for 2L MisA + CE, while still achieving sample efficiencies above $1\%$. This demonstrates that the sample efficiency remains high even for high-SNR events, highlighting that NPE may represent a promising approach to address the parameter estimation challenges posed by XG detectors.
@@
To ensure an unbiased evaluation of parameter estimation performance, we include only injections with sample efficiency exceeding $1\%$ across all seven detector configurations. The bottom panel of Figure 1 illustrates the parameter space of the retained sources, which account for 78% of the total, showing that events with $\mathcal{M}_{d}\lesssim 150\,\mathrm{M}_{\odot}$ and $D_{\mathrm{L}}\lesssim 6$ Gpc are almost entirely excluded. These sources have low sample efficiencies because their high SNR makes them more challenging for NPE to model, as their posteriors are much narrower than the prior. This behavior is further evident in the top panel of Figure 1, where sample efficiency drops sharply for sources with optimal SNR around $\sim 1000$.
@@
We are currently exploring several strategies to improve sample efficiency in regions of parameter space where performance remains below threshold, including training Dingo-IS with neural population priors (e.g., Wouters:2025csq). The results of these efforts will be presented in future work.
We further restrict the dataset to sources with optimal SNR $\rho>8$.
III.2 Multimodal posterior of a single event
@@
Figure 2: Top panel: Marginalized one-dimensional posterior distributions for the luminosity distance, recovered with Dingo-IS for an event observed with the $\Delta$ (dark blue) and 2L MisA (light blue) configurations. The results are compared with Bilby (orange). The vertical line marks the injected luminosity distance. The remaining parameters are shown in Appendix D. Middle panel: Posterior samples obtained with Dingo-IS in the $\Delta$ configuration for right ascension (ra) and declination (dec), color-coded by luminosity distance. The red cross marks the injected sky position. We also report the sky-localization area ($\Delta\Omega_{90\%}$) and comoving-volume localization ($\Delta V^{c}_{90\%}$) errors. Bottom panel: Same as the middle panel, but for the 2L MisA configuration. See Section III.2 for details.
@@
The top panel of Figure 2 shows that Dingo-IS yields posterior distributions that are statistically indistinguishable from those obtained using Bilby (see Appendix D for a formal validation). Crucially, this accuracy is achieved at a dramatically lower computational cost: Dingo-IS requires only a few minutes per source, whereas Bilby needs more than 22 hours (50 hours) to converge for this event with the $\Delta$ (2L MisA) configuration.
@@
Figure 2 also shows that Dingo-IS accurately captures the complex structure of multimodal posteriors. These multimodalities include both the eight-fold sky degeneracy arising from the triangular geometry of ET, extensively discussed in Santoliquido:2025lot, and modes introduced by the 2L MisA configuration, affecting both sky localization and luminosity distance.
@@
The bottom panel of Figure 2 shows that for the 2L MisA configuration the sky location is correlated with luminosity distance: distinct islands of probability on the sky correspond to different modes in $D_{\mathrm{L}}$. For this source, we identify three such modes. The occurrence rate of multimodal posteriors in luminosity distance is discussed in Appendix F and shown in Figure 10. The 2L MisA (2L A) configurations produce a multimodal $D_{\mathrm{L}}$ posterior more frequently than the $\Delta$ configuration, with rates of about 20% compared to $\sim$2%, respectively.
@@
These multimodalities are expected when observing short-duration, high-mass BBHs with both the 2L A and 2L MisA configurations of ET. They arise from the combined effect of the short baseline between the two interferometers ($\sim$1000 km) and the low merger frequencies of massive BBHs (typically $\lesssim 50$ Hz). We provide an illustrative example in Appendix G, showing that for a signal with higher merger frequencies, the multimodalities in both sky position and distance disappear.
@@
Despite the multimodality in $D_{\mathrm{L}}$, the number of sky-position modes is significantly reduced compared to the triangular configuration, resulting in improved sky-area and comoving-volume localization for the 2L MisA configuration.
Additional details, including the same source as observed with the 2L MisA + CE configuration, are provided in Appendix D.
III.3 Parameter estimation performance
@@
Figure 3: Cumulative distributions of events as a function of the information gain (top-left panel, see also Section II.3), the relative variations in luminosity distance (top-right panel, $\Delta D_{\mathrm{L}}/D^{\mathrm{inj}}_{\mathrm{L}}$), sky (bottom-left panel, $\Delta\Omega_{90\%}$, see also Table 1), and comoving volume localization (bottom-right panel, $V^{c}_{90\%}$) for all considered XG detector configurations (color-coded). See Section III.3 for details and Appendix H for the remaining parameters.
@@
The top-left panel of Figure 3 shows the parameter estimation performance of the considered detector configurations, measured by the information gain, revealing a clear ranking. Among the ET configurations ($\Delta$, 2L A, 2L MisA), when no other XG detectors are observing simultaneously, the 2L MisA setup performs best. Moreover, adding three LIGO observatories operating at A# sensitivity to the 2L MisA configuration further increases the information gain, underscoring the capability of this global network to estimate the parameters of the specific sources examined in this study—namely, both massive and high-redshift BBHs. The information gain further increases when 1L ET observes jointly with CE, as the longer baseline enhances both sky and volume localization. However, the separation between the 2L MisA detectors also plays a role: when observing together with CE, the 2L MisA + CE configuration outperforms $\Delta$ + CE.
@@
The top-right panel of Figure 3 shows that the luminosity distance is best measured with the $\Delta$ configuration, outperforming both the 2L A and 2L MisA configurations. Compared to Branchesi:2023mws (see their Figure 5), we find a reversed hierarchy in detector performance. We stress that our analysis is restricted to both massive and high-redshift BBH mergers. For this class of sources, the 2L configurations give rise to multimodal posteriors in $D_{\mathrm{L}}$ (see Figure 2 and Appendix F), an effect that is not captured with Fisher-matrix–based analyses.
@@
Nevertheless, the 2L MisA configuration leads to better sky and volume localization, allowing it to outperform the $\Delta$ and 2L A setup in these metrics (bottom panels of Figure 3). Improving the accuracy of the localization volume critically enhances the prospects for dark siren cosmology, which relies on the statistical association between the GW source and galaxies contained within the reconstructed localization volume (Schutz:1986gp; DelPozzo:2011vcw; Chen:2017rfc; Libanore:2020fim; Gair:2022zsa; Bosi:2023amu; Borghi:2023opd).
In Appendix H, we present the distribution of optimal SNR for individual sources, along with the performance of the detector configurations on other parameters—including the detector-frame chirp mass, mass ratio, aligned spins of both components, and inclination angle—which lead to broadly similar conclusions.
@@
Table 1: Percentage of injections (color-coded) with estimated sky localization errors within the specified areas (columns) for different detector configurations (rows). See Section III.3 for details.
@@
Figure 4: Percentage of sky maps ($y$-axis) with $\Delta\Omega_{90\%}\leq 100$ deg2 exhibiting one to eight or more disconnected modes (color-coded), for all detector configurations ($x$-axis). See Section III.3 for details.
@@
Figure 4 shows that more than 80% of events exhibit eight sky modes in the $\Delta$ configuration. This fraction is substantially reduced for the two L-shaped detector configurations: in the 2L MisA setup, the occurrence of eight sky modes drops to about 20% of the events.
The 2L MisA configuration exhibits an eight-sky-mode pattern for events lying close to the local horizon corresponding to the midpoint between the two L-shaped interferometers. For such events, the time-of-arrival difference is negligible, effectively making the two interferometers colocated (Marsat:2020rtl).
The 2L A setup localizes nearly 30% of the sources into two distinct sky modes, outperforming the 2L MisA configuration, where this fraction remains below 10%. However, the average sky-localization area per mode is significantly larger in 2L A compared to 2L MisA.
@@
None of the ET configurations are able to localize events into a single sky mode unless ET operates as part of a global network that includes either the three LIGO interferometers at A# sensitivity (including LIGO–India) or CE.
@@
The performance of ET configurations in estimating parameters of compact binary coalescences arises from the interplay between polarization measurements and a baseline that enables triangulation (Fairhurst:2009tc). Breaking the degeneracy between $D_{\mathrm{L}}$ and $\theta_{\mathrm{jn}}$ requires measuring both polarizations (Usman:2018imj), and the ability to do so is quantified by the alignment factor (Klimenko:2006rh; deSouza:2023gjv; Mascioli:2025cnw). A network composed of two L-shaped interferometers that are perfectly aligned provides no information on the cross-polarization. In contrast, the $\Delta$ configuration exhibits the highest alignment factor, yielding the most accurate polarization measurements. However, the sky maps produced by this configuration feature eight distinct sky modes owing to the colocated nature of the detectors (Singh:2020wsy; Singh:2021bwn; Santoliquido:2025lot). As a result, the 2L MisA configuration offers a trade-off, enabling polarization measurements while reducing the number of sky modes (see Figure 4).
IV Conclusions

### v2

@@ 2.3 Metrics to assess detector performance
which quantifies the extent to which the prior volume has been constrained by the data, i.e., by a factor of $2^{I}$ (Skilling, 2004). As opposed to single- or few-parameter metrics, the information gain fully characterizes the size of the 11-dimensional posterior distribution, accounting for correlations across any parameters. Equation 3 was evaluated using a Monte Carlo integral. Additional details on this are provided in Appendix E.
@@
Because Dingo-IS provides full parameter estimates for each source, we can define precise metrics to assess how different detector networks perform in source localization. For each event, we computed the area (in deg2) enclosed by the smallest 90% confidence region. The sky maps were sampled using equal-area HEALPix (Hierarchical Equal Area isoLatitude Pixelization) pixels (Gorski et al., 1999; Górski et al., 2005), where each of the $N$ pixels carries a posterior probability ($p_{i}$) that the source lies within that pixel.
@@
We employed the ligo-skymap library (Singer and Price, 2016; Singer and others, 2016a, b), which also identifies the number of disconnected sky modes (or probability islands) within a given area using a flood-fill algorithm (Foley et al., 1996), which finds all pixels connected to a starting one. Additionally, by including the posterior on luminosity distance, we computed the 90% highest probability density comoving volume ($\Delta V^{c}_{90\%}$; see Sect. 3 of Singer and others, 2016a).
@@
The other reported interval widths, for example $\Delta x/x^{\mathrm{inj}}$, indicate the relative variation of a parameter from its injected value. Here, $\Delta x$ corresponds to the square root of the associated covariance matrix element and represents the $1\sigma$ uncertainty of the parameter estimate.
3 Results
3.1 High signal-to-noise ratio versus high sample efficiency
@@
Figure 1: Top panel: Injections as a function of sample efficiency ($\epsilon$, $x$-axis) and optimal S/N ($\rho$, $y$-axis) for each detector configuration (color-coded). Filled markers indicate injections with sample efficiencies $1\%$. Percentages in parentheses denote the fraction of sources with a sample efficiency above $1\%$, while the horizontal colored lines indicate the median $\rho$ for sources exceeding this threshold. Bottom panel: Injections as a function of detector-frame chirp mass ($x$-axis) and luminosity distance (left $y$-axis), with corresponding redshifts (right $y$-axis). Colored markers indicate events with a sample efficiency greater than $1\%$ in all considered detector configurations. See Sect. 3.1 for details.
We generated $1000$ sets of BBH merger parameters sampled from the prior and used the same samples for all configurations. From the Dingo proposal distribution $q_{\phi}(\theta|d)$, we drew $10^{5}$ posterior samples and computed the corresponding importance weights (see Sect. 2.2). For each injected signal, we then evaluated the sample efficiency. To ensure unbiased estimates of the inferred parameters, we required at least $10^{3}$ effective samples per source, corresponding to a sample efficiency >1>1 %.
@@
The top panel of Fig. 1 shows that a large fraction of injected signals achieves a sample efficiency greater than $1\%$, ranging from $96\%$ for $\Delta$ and 2L A configurations to a minimum of $85\%$ for 2L MisA + CE. These values remain consistently high, with even better performance for the $\Delta$ configuration than previously reported in Santoliquido and others (2025). This improvement arises from the increased dimensionality of the embedding network (see Sect. 2.2), emphasizing the crucial role of this component within the NPE framework.
@@
The top panel of Fig. 1 also shows that $50\%$ of the events have an optimal signal-to-noise ratio (S/N; see Appendix D) exceeding $40$ for the $\Delta$ configuration and up to $70$ for 2L MisA + CE, while still achieving sample efficiencies above $1\%$. This demonstrates that the sample efficiency remains high even for high-S/N events, indicating that NPE may represent a promising approach to addressing the parameter estimation challenges posed by XG detectors.
@@
To ensure an unbiased evaluation of parameter estimation performance, we included only injections with sample efficiencies exceeding $1\%$ across all seven detector configurations. We further restricted the dataset to sources with optimal S/N $\rho>8$. The bottom panel of Fig. 1 illustrates the parameter space of the retained sources, which account for 78% of the total, showing that events with $\mathcal{M}_{d}\lesssim 150\,\mathrm{M}_{\odot}$ and $D_{\mathrm{L}}\lesssim 6$ Gpc are almost entirely excluded. These sources have low sample efficiencies because their high S/N makes them more challenging for NPE to model, as their posteriors are much narrower than the prior. This behavior is further evident in the top panel of Fig. 1, where sample efficiency drops sharply for sources with optimal S/Ns of around $\sim 1000$.
@@
We are currently exploring several strategies to improve sample efficiency in regions of parameter space where performance remains below the threshold, including training Dingo-IS with neural population priors (e.g., Wouters et al., 2025). The results of these efforts will be presented in future work.
3.2 Multimodal posterior of a single event
@@
Figure 2: Top panel: Marginalized one-dimensional posterior distributions for the luminosity distance, recovered with Dingo-IS, for an event observed with the $\Delta$ (dark blue) and 2L MisA (light blue) configurations. The results are compared with Bilby (orange). The vertical line marks the injected luminosity distance. The remaining parameters are shown in Appendix D. Middle panel: Posterior samples obtained with Dingo-IS in the $\Delta$ configuration for right ascension (ra) and declination (dec), color-coded by luminosity distance. The red cross marks the injected sky position. We also report the sky-localization area ($\Delta\Omega_{90\%}$) and comoving-volume localization ($\Delta V^{c}_{90\%}$) errors. Bottom panel: Same as the middle panel but for the 2L MisA configuration. See Sect. 3.2 for details.
@@
The top panel of Fig. 2 shows that Dingo-IS yields posterior distributions that are statistically indistinguishable from those obtained using Bilby (see Appendix D for a formal validation). Crucially, this accuracy is achieved at a dramatically lower computational cost: Dingo-IS requires only a few minutes per source, whereas Bilby needs more than 22 hours (50 hours) to converge for this event with the $\Delta$ (2L MisA) configuration.
@@
Figure 2 also shows that Dingo-IS accurately captures the complex structure of multimodal posteriors. These multimodalities include both the eight-fold sky degeneracy arising from the triangular geometry of ET, extensively discussed in Santoliquido and others (2025), and modes introduced by the 2L MisA configuration, which affect both sky localization and luminosity distance.
@@
The bottom panel of Fig. 2 shows that for the 2L MisA configuration, the sky location is correlated with luminosity distance: distinct islands of probability on the sky correspond to different modes in $D_{\mathrm{L}}$. For this source, we identify three such modes. The occurrence rate of multimodal posteriors in luminosity distance is discussed in Appendix F and shown in Fig. 11. The 2L MisA (2L A) configurations produce a multimodal $D_{\mathrm{L}}$ posterior more frequently than the $\Delta$ configuration, with rates of about 20% compared to $\sim$2%. In Appendix G we show that the multimodalities in luminosity distance and sky localization observed for the 2L configurations are robust features, persisting across different waveform approximants.
@@
These multimodalities are expected when observing short-duration, high-mass BBHs with both the 2L A and 2L MisA configurations of ET. They arise from the combined effect of the short baseline between the two interferometers ($\sim$1000 km) and the low merger frequencies of massive BBHs (typically $\lesssim 50$ Hz). We provide an illustrative example in Appendix H, showing that for a signal with higher merger frequencies, the multimodalities in both sky position and distance disappear.
@@
Despite the multimodality in $D_{\mathrm{L}}$, the number of sky-position modes is significantly reduced compared to the triangular configuration, resulting in improved sky-area and comoving-volume localization for the 2L MisA configuration. Additional details, including the same source as observed with the 2L MisA + CE configuration, are provided in Appendix D.
3.3 Parameter estimation performance
@@
Figure 3: Cumulative distributions of events as a function of the information gain (top-left panel; see also Sect. 2.3), the relative variation in luminosity distance (top-right panel; $\Delta D_{\mathrm{L}}/D^{\mathrm{inj}}_{\mathrm{L}}$), sky (bottom-left panel; $\Delta\Omega_{90\%}$; see also Table 1), and comoving volume localization (bottom-right panel; $V^{c}_{90\%}$) for all considered XG detector configurations (color-coded). See Sect. 3.3 for details.
@@
The top-left panel of Fig. 3 shows the parameter estimation performance of the considered detector configurations, measured by the information gain, revealing a clear ranking. Among the ET configurations ($\Delta$, 2L A, and 2L MisA), when no other XG detectors are observing simultaneously, the 2L MisA setup performs best. Moreover, adding three LIGO observatories operating at A# sensitivity to the 2L MisA configuration further increases the information gain, underscoring the capability of this global network to estimate the parameters of the specific sources examined in this study—namely, both massive and high-redshift BBHs. The information gain further increases when 1L ET observes jointly with CE, as the longer baseline enhances both sky and volume localization. However, the separation between the 2L MisA detectors also plays a role: when observing together with CE, the 2L MisA + CE configuration outperforms $\Delta$ + CE.
@@
The top-right panel of Fig. 3 shows that the luminosity distance is best measured with the $\Delta$ configuration, outperforming both the 2L A and 2L MisA configurations. Compared to Branchesi and others (2023, see their Fig. 5), we find a reversed hierarchy in detector performance. We stress that our analysis is restricted to both massive and high-redshift BBH mergers. For this class of source, the 2L configurations give rise to multimodal posteriors in $D_{\mathrm{L}}$ (see Fig. 2 and Appendix F), an effect that is not captured with Fisher-matrix-based analyses.
@@
Nevertheless, the 2L MisA configuration leads to better sky and volume localization, allowing it to outperform the $\Delta$ and 2L A setup in these metrics (bottom panels of Fig. 3). Improving the accuracy of the localization volume critically enhances the prospects for dark siren cosmology, which relies on the statistical association between the GW source and galaxies contained within the reconstructed localization volume (Schutz, 1986; Del Pozzo, 2012; Chen et al., 2018; Libanore et al., 2021; Gair and others, 2023; Bosi et al., 2023; Borghi et al., 2024).
Table 1: Percentage of injections (color-coded) with estimated sky localization errors within the specified areas (columns) for different detector configurations (rows). [^1] See Sect. 3.3 for details.
@@
Figure 4: Percentage of sky maps ($y$-axis) with $\Delta\Omega_{90\%}\leq 100$ deg2 exhibiting one to eight or more disconnected modes (color-coded), for all detector configurations ($x$-axis). See Sect. 3.3 for details.
@@
Figure 4 shows that more than 80% of events exhibit eight sky modes in the $\Delta$ configuration. This fraction is substantially reduced for the two L-shaped detector configurations: in the 2L MisA setup, the occurrence of eight sky modes drops to about 20% of the events. The 2L MisA configuration exhibits an eight-sky-mode pattern for events located close to the local horizon defined at the midpoint between the two L-shaped interferometers. For such events, the time-of-arrival difference is negligible, effectively making the two interferometers colocated (Marsat et al., 2021).
@@
The 2L A setup localizes nearly 30% of the sources in two distinct sky modes, outperforming the 2L MisA configuration, for which this fraction remains below 10%. However, the average sky-localization area per mode is significantly larger in 2L A compared to 2L MisA. None of the ET configurations are able to localize events in a single sky mode unless ET operates as part of a global network that includes either the three LIGO interferometers at A# sensitivity (including LIGO–India) or CE.
For completeness, Fig. 5 presents the performance of the considered detector configurations in estimating the remaining parameters, such as optimal S/N, detector-frame chirp mass, mass ratio, the aligned spin components, and inclination angle. The overall trend is confirmed, as the $\Delta$ and 2L A configurations provide the poorest performance, while the 2L MisA + CE setup delivers the most accurate estimates.
@@
The performance of ET configurations in estimating parameters of compact binary coalescences arises from the interplay between polarization measurements and a baseline that enables triangulation (Fairhurst, 2009). Breaking the degeneracy between $D_{\mathrm{L}}$ and $\theta_{\mathrm{jn}}$ requires measuring both polarizations (Usman et al., 2019), and the ability to do so is quantified by the alignment factor (Klimenko et al., 2006; de Souza and Sturani, 2023; Mascioli et al., 2025). A network composed of two L-shaped interferometers that are perfectly aligned provides no information on the cross-polarization. In contrast, the $\Delta$ configuration exhibits the highest alignment factor, yielding the most accurate polarization measurements. However, the sky maps produced by this configuration feature eight distinct sky modes owing to the colocated nature of the detectors (Singh and Bulik, 2021, 2022; Santoliquido and others, 2025). As a result, the 2L MisA configuration offers a tradeoff, enabling polarization measurements while reducing the number of sky modes (see Fig. 4).
Figure 5: Cumulative distributions of events as a function of the optimal S/N ($\rho$; Eq. 11), the relative variations in detector-frame chirp mass ($\Delta\mathcal{M}_{\mathrm{d}}/\mathcal{M}^{\mathrm{inj}}_{\mathrm{d}}$), mass ratio ($q$), first ($\Delta\chi_{1}$) and second ($\Delta\chi_{2}$) aligned spins, and inclination angle ($\Delta\theta_{\mathrm{jn}}$) for all considered XG detector configurations. See Sect. 3.3 for details.
4 Conclusions